You have two cups, one with m1 = 0.92 kg of water at t1 = 29.76oC and the other with m2 = 2.88 kg at t2 = 45.6oC. You mix the two in an insulated thermos. What is the final temperature of the water?
Answer #1

Heat gained= Heat lost

m1*s*(T-t1)=m2*s*(t2-T)

0.92*(T-29.76)=2.88*(45.6-T)

T=41.77in degree celsius